OpenStudy (anonymous):

What is the lateral area of a cylinder with a radius of 2 in and a height of 5 in? Use 3.14 for

OpenStudy (kropot72):

The circumference of the cylinder is 3.14 * (2 * 2) in = 12.56 in Do you follow?

OpenStudy (kropot72):

The above comes from the formula for circumference: Circumference = \[\pi \times d\] where d is the diameter of a circle. In this case the diameter is the radius multiplied by two.

OpenStudy (kropot72):

The lateral area is found by multiplying the circumference by the height. Can you do that and post your result?

OpenStudy (anonymous):

39.4

OpenStudy (kropot72):

The circumference is 12.56 inches and the height is 5 inches. 12.56 * 5 = 62.8 square inches